OpenStudy (anonymous):

Solve 8(cos pi/2 + i sin pi/2) times (1/2(cos pi/6 + i sin pi/6)

OpenStudy (sirm3d):

multiplication of complex numbers in trigonometric form follow this rule: \[\left[ r_1(\cos A + i \sin A) \right]\left[ r_2(\cos B + i \sin B)\right]=r_1r_2 \left[ \cos (A+B) + i \sin (A + B)\right]\]